# Make Me Happier: Evoking Emotions Through Image Diffusion Models

### v1

@@ Abstract
Despite the rapid progress in image generation, emotional image editing remains under-explored. The semantics, context, and structure of an image can evoke emotional responses, making emotional image editing techniques valuable for various real-world applications, including treatment of psychological disorders, commercialization of products, and artistic design. For the first time, we present a novel challenge of emotion-evoked image generation, aiming to synthesize images that evoke target emotions while retaining the semantics and structures of the original scenes. To address this challenge, we propose a diffusion model capable of effectively understanding and editing source images to convey desired emotions and sentiments. Moreover, due to the lack of emotion editing datasets, we provide a unique dataset consisting of 340,000 pairs of images and their emotion annotations. Furthermore, we conduct human psychophysics experiments and introduce four new evaluation metrics to systematically benchmark all the methods. Experimental results demonstrate that our method surpasses all competitive baselines. Our diffusion model is capable of identifying emotional cues from original images, editing images that elicit desired emotions, and meanwhile, preserving the semantic structure of the original images. All code, model, and data will be made public.
@@ 1 Introduction
What we perceive not only entails useful visual information but also evokes profound emotional reactions. Incorporating visual cues that elicit emotions into images holds practical significance in various domains. For instance, in commercials, it amplifies product branding and shapes public opinions, captivating consumer attention. Similarly, integrating emotion-evoking elements into AR/VR experiences aids in addressing psychological disorders such as autism and schizophrenia. Despite significant advancements in image generation techniques [6, 27, 35], there has been a notable absence of focus given to emotion conditioning. Here, we introduce the novel and important problem of emotion-evoked image generation — given an initial image and a target emotion, the task is to create an image that elicits the specified emotion of human viewers while maintaining the original scene’s semantics and structures.
Emotion-evoked image generation poses a challenge as it requires a deep understanding and identification of the subtle context and semantic elements present in the source images that evoke emotional responses. Additionally, this task also requires image editing capabilities to effectively remove, replace, add, or modify elements within the source images to elicit the intended emotional response. For instance, in Figure 1, the presence of a burning forest significantly contributes to feelings of anger. Consequently, in the generated image, the burning forest is substituted with a serene scene featuring a cluster of flowers in the forest, thereby transforming the overall emotional response to awe.
@@
Studies [31, 32] have identified both global and local factors influencing emotional states, such as overall color and tonal shifts in image backgrounds, local facial expressions, and the presence of emotion-associated objects like graveyards correlated with sadness and balloons correlated with amusement. As also seen in Figure 2, introducing nauseating or fearful elements like fire or ghosts into forests (Column 2) can readily evoke negative emotions, as these elements prompt instinctual responses to unpleasant or adverse environments. However, not all concepts serve as emotional triggers; some may remain neutral, such as the sky, plants, or windows. In the same examples in Column 2, forests are neutral concepts and do not elicit emotional responses. Given that human emotions tend to exhibit greater sensitivity to negative stimuli over positive stimuli [2, 11], simply manipulating global intensity by adjusting the brightness of entire images in Column 2 would not necessarily alter the negative emotions.
Drawing inspiration from these emotion-associated factors above, we propose our model, EmoEditor, for emotion-evoked image generation. We base EmoEditor on state-of-the-art stable diffusion models and introduce three technical novelties. (1) EmoEditor employs a novel dual-branch architecture in the reverse diffusion process, that integrates emotion-conditioned global context and local emotional cues from the source images, to ensure coherence with the original content while reflecting the desired emotional outcome. (2) During training, the network’s behaviors are guided by aligning model creativity with human expectations in the neuro-symbolic space, implicitly learning the mapping from target emotions to human-annotated text instructions for image editing. (3) During inference, iterative emotion discrimination mechanisms are employed to autonomously select emotionally coherent images.
Before the introduction of our EmoEditor, traditional methods, such as color transfer and style transfer [21, 7, 30], were commonly employed to modify the emotional tone of images. These methods typically require reference emotional images or textual cues as additional inputs. However, they face challenges in achieving satisfactory performance in eliciting desired emotions because they struggle to understand, identify, and manipulate the local regions of source images that evoke emotional responses. In Figure 2, while darkening brightness or deepening hues may easily evoke fear (Example 1), simply brightening colors in an image of a forest fire is inadequate for eliciting amusement (Example 4).
With the rapid advancements in diffusion models, various effective text-to-image editing techniques have emerged [12, 3, 24]. However, these methods are not specifically tailored for understanding and identifying emotion-associated elements within images. To address this gap, we propose a competitive baseline (Large Model Series), concatenating multiple large language and vision models. This baseline first understands source images through image captioning, then generates emotion-prompted instructions using language models, and finally edits images prompted by these instructions with diffusion models. Upon comparison with existing methods through a series of human psychophysics experiments and four new quantitative evaluations, our end-to-end trained EmoEditor achieves outstanding image generation results. It effectively preserves the structural coherence and semantic consistency wtih the source images while eliciting the target emotion of human viewers.
To obtain training data for emotion-evoked image generation, we establish a dataset curation pipeline and introduce the first large-scale EmoPair dataset, comprising approximately 340,000 pairs of images annotated with emotion-conditioned editing instructions. Unlike existing datasets on visual emotion analysis [18, 17, 34, 20, 31], which assign emotion labels to individual images for classification, our dataset provides paired images annotated with source and target emotion labels.
@@ 2.1 Visual Emotion Analysis
Visual Emotion Analysis (VEA) aims to understand and predict human emotional responses through visual information from images or videos. In psychology, Categorical Emotion States (CES) form the basis for emotion classification models, notably exemplified by the 6-category Ekman model [5] and the 8-category Mikels model [18]. Early research focused on manually crafted features such as color and content [17], while recent advancements leverage deep learning models to automatically capture intrinsic global and local emotional cues for more accurate predictions [33, 32]. Despite advancements in visual emotion analysis, there is a limited focus on generating images that evoke specific emotional responses. To address this, we propose EmoEditor, a two-branch diffusion model that integrates emotion recognition and image generation in a unified framework. Unlike Affective Image Filter (AIF) [30] which relies on explicit textual instructions, EmoEditor does not need explicit text guidance on scene modifications for desired emotions.
Exsiting VEA datasets [20, 17, 34] primarily focus on emotion classification, including EmoSet [31], the largest dataset with 120,000 images annotated for emotion labels and six attributes. However, none of them include image pairs with annotated emotion labels and text editing instructions crucial for emotion-evoked image generation. Therefore, we introduce the EmoPair dataset, comprising 340,000 image pairs with differing source and target emotions while ensuring consistent scene semantics and structure.
@@ 2.2 Image Editing
Color and Style Transfer. Color and style transfer methods like PDF-Transfer[21], Neural Style Transfer[7], and CLIP-Styler[14] use reference images for image manipulation, with outcomes varying based on choices. While colors and styles may influence human emotions, they struggle to manipulate local image regions, limiting their versatility in emotion-evoked generation. In contrast, our EmoEditor balances local and global image manipulations during inference without requiring additional references.
Text-to-Image Editing. Text-to-image editing is a growing area of interest. Previous studies utilize Generative Adversarial Networks (GAN) [8] architectures and leverage CLIP model’s [22] language embedding for image manipulation from detailed text descriptions[19, 1]. Recently, text-to-image diffusion models like Imagic[12], Prompt-to-prompt[10], InstructPix2Pix[3], and DreamBooth[24] have strong capabilities in generating high-quality and diverse images based on text prompts. Despite their impressive results, they often overlook emotional aspects and rely heavily on detailed textual instructions. In contrast, our EmoEditor only needs target emotion and source image inputs. Through end-to-end training, it comprehends emotional cues and creatively edits images for the desired emotions while preserving structure and coherence.
@@ 3 Method
Figure 3: Pipeline for Curating Our EmoPair Dataset. The dataset comprises two subsets: EmoPair-Annotated Subset (EPAS, left blue box) and EmoPair-Generated Subset (EPGS, right orange box). Each subset includes schematics depicting the creation, selection, and labeling of image pairs in the upper quadrants, with two example pairs in the lower quadrants. Each example pair comprises a source image (framed in green) and a target image. The classified source and target emotion labels (highlighted in red) and target-emotion-driven text instructions for image editing are provided. See more examples in Supp.. See Sec. 3.1 for EmoPair dataset.
@@ 3.1 Curating Our EmoPair Dataset
Given the unavailability of a paired image dataset for emotion editing, we curate EmoPair. It consists of two subsets: EmoPair-Annotated Subset (EPAS), containing 331,595 image pairs from Ip2p[3] annotated with emotion labels; and EmoPair-Generated Subset (EPGS), consisting of 6,949 pairs generated based on text instructions given target emotions. All images from EmoPair are of size 224$\times$224. Similar to existing works in VEA [18, 31], all images are classified into 8 emotion categories, encompassing four positive (amusement, awe, contentment, excitement) and four negative (anger, disgust, fear, sadness) emotions. Figure 3 shows the pipeline for curating our EmoPair dataset.
@@
EmoPair-Annotated Subset (EPAS). The original Ip2p dataset [3] lacks emotion labels, consisting only of image pairs with text instructions for editing. To address this limitation, we use an emotion predictor $\mathcal{P}$ to classify the source and target images from the Ip2p dataset into eight distinct emotions. $\mathcal{P}$ employs ResNet18 [9] as a backbone and trained on the EmoSet dataset [31]. It achieves a top-1 accuracy of 73% on EmoSet’s test set. We augment the Ip2p dataset with emotion labels, resulting in a total of 331,595 image pairs. Notably, the text instructions with these image pairs serve as human expectations to guide EmoEditor’s training (Sec. 3.2), but are not utilized during inference.
EmoPair-Generated Subset (EPGS). Due to the lack of salient emotional cues in the Ip2p dataset, we further complement EPAS with EPGS subset, which has emotionally rich images from the EmoSet dataset [31]. EmoSet provides diverse and evocative emotional cues that benefit our EmoEditor. However, EmoSet lacks paired images with coherent scene structures and target emotions, along with the absence of text editing instructions from source to target images.
@@
Directly inputting a single-word prompt like “[Target Emotion]" into existing text-to-image editing models proves ineffective due to their limited emotional understanding and reasoning capabilities. Single-emotion terms lack specificity, resulting in varied interpretations depending on the context of source images. For instance, enhancing an image for excitement can vary significantly based on context, such as prompting a happy smile for a person or adding fireworks to the sky for a landscape. Thus, to ensure accurate image pair generation, we created 50 general instructions, agnostic of source images, to prompt transitions to desired emotions across 8 categories using GPT-3 [4]. Human annotators then ranked these instructions per emotion category based on efficacy and retained the top ten. Next, the pre-trained Ip2p model manipulated EmoSet source images using these selected instructions to generate target images.
We employ the following selection criteria to control the quality of generated image pairs: (1) Using our emotion predictor $\mathcal{P}$, we analyze the generated images and only select those with a Top-1 classification confidence over 90% for the target emotions. (2) To ensure the preservation of similar scene structures, we utilize Structural Similarity Index (SSIM) [29] and Learned Perceptual Image Patch Similarity (LPIPS) [36] for filtering the remaining outcomes. SSIM measures structural similarity between images, while LPIPS quantifies perceptual differences. Specifically, we require the generated images $\hat{x}$ and the source image $x$ to meet the conditions below: 0.3 < SSIM($x$, $\hat{x}$) < 0.6 and LPIPS($x$, $\hat{x}$) > 0.1. Ultimately, the EPGS subset retains 6,949 image pairs.
@@ 3.2 Our Diffusion Model - EmoEditor
Our EmoEditor, in Figure 4, takes a source image $I_{s}$ and a target emotion $y$ to generate an image evoking the desired emotion. It employs the latent diffusion model (LDM) [23] with three key novelties. EmoEditor avoids using emotion labels for $I_{s}$ for two reasons: (1) Obtaining emotion labels for source images is often impractical in real-world scenarios. (2) Source images can be complex, conveying multiple emotions simultaneously. Utilizing a stereotypical source emotion as input potentially introduces biases and limits the model’s ability to generate diverse outputs.
Integration of Global and Local Emotion Cues. In the global branch of our EmoEditor, we convert the target emotion $y$ into a binary one-hot vector $e_{oh}$, where the corresponding entry to $y$ in $e_{oh}$ is set to 1 and the rest are set to 0. Next, we introduce an emotion encoder $\tau_{\theta}$, which encodes $e_{oh}$ to $e$ via a stack of fully connected blocks: $e=\tau_{\theta}(e_{oh})$.
@@
Unlike classical text-to-image models using explicit text prompts, our $\tau_{\theta}$ is trained from scratch, learning neuro-symbolic embeddings for the reverse diffusion process without relying on pre-trained language models like CLIP[22]. Later on, we introduce novel losses to regularize the behaviors of $\tau_{\theta}$. Notably, we intentionally avoid using explicit text editing instructions from image pairs as conditional inputs to the model due to the varied interpretations of target emotions, where multiple solutions can elicit the same emotions. Including explicit text instructions may hinder the model’s creativity, restricting its capacity to produce diverse sets of images conveying the same target emotions.
@@
In the local branch of our EmoEditor, we introduce an image encoder $\mathcal{E}$ to extract visual features of the source image, acquiring $z$ for $I_{s}$: $z=\mathcal{E}(I_{s})$. Similar to the LDM [23], $\mathcal{E}$ is based on a pre-trained variational autoencoder[13]. The forward diffusion process gradually adds noise to $z$, generating a noisy latent $z_{t}$, where the noise level increases over $t$ time steps and $t\in T$. Specifically, we expressed $z_{t}$ as $z_{t}=\sqrt{\alpha_{t}}\cdot z_{t-1}+\sqrt{1-\alpha_{t}}\cdot\epsilon$, where $\alpha_{t}$ is the diffusion coefficient controlling the rate of noise increase, and $\epsilon$ is a random noise sampled from a normal distribution with mean 0 and variance 1.
@@
In the reverse diffusion process, the denoising network $\epsilon_{\theta}$ predicts noise added to the noisy latent $z_{t}$ using cross-attention with the target emotion condition $e$ and source image condition $z$. The features in $z$ and $z_{t}$ enhance understanding of intrinsic information in the image, leading to emotion-specific content for $I_{s}$. After $T-1$ denoising steps, decoder $\mathcal{D}$ produces an emotion-evoked image.
We minimize the latent diffusion loss, which is the expected error between the predicted noise by $\epsilon_{\theta}$ and the actual noise $\epsilon$ sampled during training: $L_{noise}=E_{t,y,z,\epsilon(z_{t}),\epsilon\sim N(0,1)}\left\|\epsilon-\epsilon_{\theta}(z_{t},t,z,\tau_{\theta}(y))\right\|^{2}$.
@@
Alignment Loss between Model and Human Judgements. A single emotion term lacks the depth to convey the nuanced causes underlying emotions, posing a challenge for diffusion models to grasp the diagnostic information triggering emotions. In contrast to existing text-to-image models, EmoEditor emulates human reasoning by aligning its thought process with text instructions in our EmoPair dataset, which leads the transition from the source emotion to the target emotion. To achieve this, we introduce an alignment loss in a neuro-symbolic space, between emotion embeddings $e$ from $\tau_{\theta}$ and text embeddings $c$ of instructions in each image pair of our EmoPair dataset with pre-trained CLIP [22]. Specifically, the alignment loss is formulated as the inverse of cosine similarity between $e$ and $c$: $L_{emb}=1-\cos(e,c)$. Overall, we conduct joint training of $\tau_{\theta}$ and $\epsilon_{\theta}$ with the following loss: $L_{total}=L_{noise}+\lambda L_{emb}$, where $\lambda=0.5$ represents the weight for balancing different losses.
@@
In the first iteration of image generation, the source image $I_{s}$ serves as both the input and condition image, producing latent variables $z$ and $z_{t}$. These, along with the emotion vector $e$, are fed to the EmoEditor to obtain the generated image $\hat{I}$. The critic evaluates the generated image quality based on two criteria: (1) the structural similarity (SSIM) between $\hat{I}$ and $I_{s}$ falls within 0.3-0.8, and (2) the predicted emotion by $\mathcal{P}$ on $\hat{I}$ matches the target emotion, with a confidence level exceeding 0.8. If both criteria are met, EmoEditor ceases image generation; otherwise, it continues to generate new images. In subsequent iterations, we use $\hat{I}$ as the input image, generating its noisy latent variable $\hat{z}_{t}$, with $I_{s}$ remaining as the condition. Along with the emotion vector $e$, the EmoEditor generates new edited images for the critic. This process iterates until both criteria are satisfied, or the number of critic iterations exceeds a predefined limit of 20. In the latter case, we select the generated result with the highest confidence level predicted by $\mathcal{P}$ given the target emotion among all past iterations.
Implementation Details. Follow Wang et al.[28], we initialize the weights of $\mathcal{E}$, $\mathcal{D}$, and $\epsilon_{\theta}$ with the pre-trained Ip2p[3] weights. Throughout the training process, we maintain the fixed parameters of $\mathcal{E}$ and $\mathcal{D}$, focusing on training $\tau_{\theta}$ and $\epsilon_{\theta}$. The frozen emotion predictor $\mathcal{P}$ is only used during inference. We conduct experiments with NVIDIA RTX A5000 GPUs, implementing our PyTorch-based framework using the ADAM optimizer. The max time step $T$ is set to 1,000.

### v2

@@ Abstract
Despite the rapid progress in image generation, emotional image editing remains under-explored. The semantics, context, and structure of an image can evoke emotional responses, making emotional image editing techniques valuable for various real-world applications, including treatment of psychological disorders, commercialization of products, and artistic design. For the first time, we present a novel challenge of emotion-evoked image generation, aiming to synthesize images that evoke target emotions while retaining the semantics and structures of the original scenes. To address this challenge, we propose a diffusion model capable of effectively understanding and editing source images to convey desired emotions and sentiments. Moreover, due to the lack of emotion editing datasets, we provide a unique dataset consisting of 340,000 pairs of images and their emotion annotations. Furthermore, we conduct human psychophysics experiments and introduce four new evaluation metrics to systematically benchmark all the methods. Experimental results demonstrate that our method surpasses all competitive baselines. Our diffusion model is capable of identifying emotional cues from original images, editing images that elicit desired emotions, and meanwhile, preserving the semantic structure of the original images. All code, model, and dataset will be made public.
@@ 1 Introduction
What we perceive not only entails useful visual information but also evokes profound emotional reactions. Incorporating visual cues that elicit emotions into images holds practical significance in various domains. For instance, in commercials, it amplifies product branding and shapes public opinions, captivating consumer attention. Similarly, integrating emotion-evoking elements into AR/VR experiences aids in addressing psychological disorders such as autism and schizophrenia. Despite significant advancements in image generation techniques Gal et al. (2022); Wang et al. (2023); Yu et al. (2018), there has been a notable absence of focus given to emotion conditioning. Here, we introduce the novel and important problem of emotion-evoked image generation — given an initial image and a target emotion, the task is to create an image that elicits the specified emotion of human viewers while maintaining the original scene’s semantics and structures.
Emotion-evoked image generation poses a challenge as it requires a deep understanding and identification of the subtle context and semantic elements present in the source images that evoke emotional responses. Additionally, this task also requires image editing capabilities to effectively remove, replace, add, or modify elements within the source images to elicit the intended emotional response. For instance, in Figure 1, the presence of a burning lawn significantly contributes to feelings of anger. Consequently, in the generated image, the burning lawn is substituted with a serene scene featuring a cluster of flowers, thereby transforming the overall emotional response to awe.
Studies Yang et al. (2023, 2018) have identified both global and local factors influencing emotional states, such as overall color and tonal shifts in image backgrounds, local facial expressions, and the presence of emotion-associated objects like graveyards correlated with sadness and balloons correlated with amusement. As also seen in Figure 2, introducing fearful elements like fire or a woman in white into forests (Column 2) can readily evoke negative emotions, as these elements prompt instinctual responses to unpleasant or adverse environments. However, not all concepts serve as emotional triggers; some may remain neutral, such as the sky, plants, or windows. In the same examples in Column 2, forests are neutral concepts and do not elicit emotional responses. Given that human emotions tend to exhibit greater sensitivity to negative stimuli over positive stimuli Baumeister et al. (2001); Ito et al. (1998), simply manipulating global intensity by adjusting the brightness of entire images in Column 2 would not necessarily alter the negative emotions.
Drawing inspiration from these emotion-associated factors, we propose our model, EmoEditor, for emotion-evoked image generation. We base EmoEditor on state-of-the-art stable diffusion models and introduce three technical novelties. (1) EmoEditor employs a novel dual-branch architecture in the reverse diffusion process, that integrates emotion-conditioned global context and local emotional cues from the source images, to ensure coherence with the original content while reflecting the desired emotional outcome. (2) During training, the network’s behaviors are guided by aligning model creativity with human expectations in the neuro-symbolic space, implicitly learning the mapping from target emotions to human-annotated text instructions. (3) During inference, iterative emotion discrimination mechanisms are employed to autonomously select emotionally coherent images.
Before the introduction of our EmoEditor, traditional methods, such as color transfer and style transfer Pitié et al. (2007); Gatys et al. (2015); Weng et al. (2023), were commonly employed to modify the emotional tone of images. These methods typically require reference emotional images or textual cues as additional inputs. However, they face challenges in achieving satisfactory performance in eliciting desired emotions because they struggle to understand, identify, and manipulate the local regions of source images that evoke emotional responses. In Figure 2, while darkening brightness or deepening hues may easily evoke fear (Example 1), simply brightening colors in an image of a forest fire is inadequate for eliciting amusement (Example 4).
@@
With the rapid advancements in diffusion models, various effective text-to-image editing techniques have emerged Kawar et al. (2023); Brooks et al. (2023); Ruiz et al. (2023). However, these methods are not specifically tailored for understanding and identifying emotion-associated elements within images. To address this gap, we propose a competitive baseline (Large Model Series), concatenating multiple large language and vision models. This baseline first understands source images through image captioning, then generates emotion-prompted instructions using language models, and finally edits images prompted by these instructions with diffusion models. Upon comparison with existing methods through a series of human psychophysics experiments and four new quantitative evaluations, our end-to-end trained EmoEditor achieves outstanding image generation results. It effectively preserves the structural coherence and semantic consistency wtih the source images while eliciting the target emotion of human viewers.
To obtain training data for emotion-evoked image generation, we establish a dataset curation pipeline and introduce the first large-scale EmoPair dataset, comprising approximately 340,000 pairs of images annotated with emotion-conditioned editing instructions. Unlike existing datasets on visual emotion analysis Mikels et al. (2005); Machajdik, Hanbury (2010); You et al. (2016); Peng et al. (2015); Yang et al. (2023), which assign emotion labels to individual images for classification, our dataset provides paired images annotated with source and target emotion labels.
@@ 2.1 Visual Emotion Analysis
Visual Emotion Analysis (VEA) aims to understand and predict human emotional responses from visual data. In psychology, Categorical Emotion States (CES) form the basis for emotion classification models like the 8-category Mikels model Mikels et al. (2005). Early research used manually crafted features such as color and content Machajdik, Hanbury (2010), while recent advancements employ deep learning to capture emotional cues more accurately You et al. (2015); Yang et al. (2018). Despite progress, generating images to evoke specific emotions remains underexplored. We propose EmoEditor, a two-branch diffusion model that integrates emotion recognition and image generation to address this. Unlike the Affective Image Filter Weng et al. (2023), which requires explicit textual instructions, EmoEditor does not need text guidance to achieve desired emotions.
Existing VEA datasets Peng et al. (2015); You et al. (2016) focus on emotion classification, including EmoSet Yang et al. (2023), which has 120k images with emotion labels and attributes. However, they lack image pairs with annotated emotions and text editing instructions essential for emotion-evoked image generation. Thus, we introduce the EmoPair dataset, comprising 340k image pairs with differing source and target emotions while ensuring consistent scene semantics and structure.
@@ 2.2 Image Editing
Color and Style Transfer. Methods like PDF-TransferPitié et al. (2007), Neural Style TransferGatys et al. (2015), and CLIP-StylerKwon, Ye (2022) manipulate images using reference images, but they struggle with local region adjustments, limiting their versatility in generating emotion-evoking images. Our EmoEditor balances local and global manipulations without needing additional references.
Text-to-Image Editing. Text-to-image editing is increasingly popular. Earlier studies used GANs Goodfellow et al. (2014) and CLIP models Radford et al. (2021) for image manipulation based on detailed text descriptions Patashnik et al. (2021); Abdal et al. (2022). Recent diffusion models like Imagic Kawar et al. (2023), Prompt-to-prompt Hertz et al. (2022), InstructPix2Pix Brooks et al. (2023), and DreamBooth Ruiz et al. (2023) generate high-quality images from text prompts but often ignore emotional aspects and rely heavily on detailed instructions. In contrast, our EmoEditor needs only target emotion and source image inputs, using end-to-end training to understand and edit emotional cues while preserving image structure and coherence.
@@ 3 Method
Figure 3: Pipeline for Curating Our EmoPair Dataset. The dataset comprises two subsets: EmoPair-Annotated Subset (EPAS, left blue box) and EmoPair-Generated Subset (EPGS, right orange box). Each subset includes schematics depicting the creation, selection, and labeling of image pairs in the upper quadrants, with two example pairs in the lower quadrants. Each example pair comprises a source image (framed in green) and a target image. The classified source and target emotion labels (highlighted in red) and target-emotion-driven text instructions for image editing are provided. See more examples in Supp.Sec.S1. See Sec.3.1 for EmoPair dataset.
@@ 3.1 Curating Our EmoPair Dataset
Given the unavailability of a paired image dataset for emotion editing, we curate EmoPair. It consists of two subsets: EmoPair-Annotated Subset (EPAS), containing 331,595 image pairs from Ip2pBrooks et al. (2023) annotated with emotion labels; and EmoPair-Generated Subset (EPGS), consisting of 6,949 pairs generated based on text instructions given target emotions. All images from EmoPair are of size 224$\times$224. Figure 3 shows the pipeline for curating our EmoPair dataset.
Emotion Predictor $\mathcal{P}$. To better construct our dataset, we first train an emotion predictor $\mathcal{P}$ on the existing VEA dataset EmoSet Yang et al. (2023). All images are classified into 8 emotion categories, encompassing four positive (amusement, awe, contentment, excitement) and four negative (anger, disgust, fear, sadness) emotions. $\mathcal{P}$ employs ResNet18 He et al. (2016) as a backbone and achieves a top-1 accuracy of 73% on EmoSet’s test set.
@@
EmoPair-Annotated Subset (EPAS). The original Ip2p dataset lacks emotion labels, consisting only of image pairs with text instructions for editing. To address this, we use $\mathcal{P}$ to classify the source and target images from the Ip2p dataset into eight distinct emotions. We augment the Ip2p dataset with emotion labels, resulting in 331,595 image pairs. Notably, the text instructions with these image pairs serve as human expectations to guide EmoEditor’s training but are not utilized during inference.
@@
EmoPair-Generated Subset (EPGS). To address the lack of emotional cues in the Ip2p dataset, we create the EPGS subset using emotionally rich images from EmoSet Yang et al. (2023). However, EmoSet lacks paired images with consistent scenes and target emotions, and it doesn’t include text editing instructions. Single-word prompts like “[Target Emotion]" prove ineffective in existing text-to-image models due to their limited emotional reasoning. Single-emotion terms lack specificity, resulting in varied interpretations depending on the context of source images. For instance, enhancing an image for excitement can vary significantly based on context, such as prompting a happy smile for a person or adding fireworks to the sky for a landscape. Thus, to generate accurate image pairs, we develop 50 general instructions for transitioning to desired emotions of 8 categories using GPT-3 Brown et al. (2020). Human annotators rank these instructions, keeping the top ten per emotion. The pre-trained Ip2p model then uses these instructions to manipulate EmoSet source images, resulting in 6,949 high-quality image pairs after quality control measures (see Supp.Sec.S1).
@@ 3.2 Our Diffusion Model - EmoEditor
Our EmoEditor, in Figure 4, takes a source image $I_{s}$ and a target emotion $y$ to generate an image evoking the desired emotion. It employs the latent diffusion model (LDM) Rombach et al. (2022) with three key novelties. EmoEditor avoids using emotion labels for $I_{s}$ for two reasons: (1) Obtaining emotion labels for source images is often impractical in real-world scenarios. (2) Source images can be complex, conveying multiple emotions simultaneously. Utilizing a stereotypical source emotion as input potentially introduces biases and limits the model’s ability to generate diverse outputs.
@@
Integration of Global and Local Emotion Cues. In the global branch of our EmoEditor, we convert the target emotion $y$ into a binary one-hot vector $e_{oh}$, where the corresponding entry to $y$ in $e_{oh}$ is set to 1 and the rest are set to 0. This simplifies model design, enhancing comprehension and generation of complex emotional images, as one-hot encoding can capture intricate emotion combinations through a multi-label representation. Conversely, text inputs demand extra transformation steps, potentially introducing noise or errors. We introduce an emotion encoder $\tau_{\theta}$ to encode $e_{oh}$ into $e$ via fully connected layers: $e=\tau_{\theta}(e_{oh})$ (see Supp.Sec.S2.1 for details), trained from scratch to learn neuro-symbolic embeddings. Later on, we introduce novel losses to regularize the behaviors of $\tau_{\theta}$. Notably, we intentionally avoid using explicit text editing instructions as conditional inputs to the model due to the varied interpretations of target emotions, where multiple solutions can elicit the same emotions. Including explicit text instructions may hinder the model’s creativity, restricting its capacity to produce diverse sets of images conveying the same target emotions.
@@
In the local branch of our EmoEditor, we introduce an image encoder $\mathcal{E}$ to extract visual features of the source image, acquiring $z$ for $I_{s}$: $z=\mathcal{E}(I_{s})$. Similar to the LDM Rombach et al. (2022), $\mathcal{E}$ is based on a pre-trained VAEKingma, Welling (2022). The forward diffusion process gradually adds noise to $z$, generating a noisy latent $z_{t}$, where the noise level increases over $t$ time steps ($t\in T$). Specifically, we expressed $z_{t}=\sqrt{\alpha_{t}}\cdot z_{t-1}+\sqrt{1-\alpha_{t}}\cdot\epsilon$, where $\alpha_{t}$ is the diffusion coefficient controlling the rate of noise increase, and $\epsilon$ is a random noise sampled from a normal distribution.
@@
In the reverse diffusion process, the denoising network $\epsilon_{\theta}$ predicts noise added to the noisy latent $z_{t}$ using cross-attention with the target emotion condition $e$ and source image condition $z$. The features in $z$ and $z_{t}$ enhance understanding of intrinsic information in the image, leading to emotion-specific content for $I_{s}$. After $T-1$ denoising steps, decoder $\mathcal{D}$ produces an emotion-evoked image. We minimize the latent diffusion loss, which is the expected error between the predicted noise by $\epsilon_{\theta}$ and the actual noise $\epsilon$ sampled during training: $L_{noise}=E_{t,y,z,\epsilon(z_{t}),\epsilon\sim N(0,1)}\left\|\epsilon-\epsilon_{\theta}(z_{t},t,z,\tau_{\theta}(y))\right\|^{2}$.
@@
Alignment Loss between Model and Human Judgements. A single emotion term lacks the depth to convey the nuanced causes underlying emotions, posing a challenge for diffusion models to grasp the diagnostic information triggering emotions. In contrast to existing text-to-image models, EmoEditor emulates human reasoning by aligning its thought process with text instructions in our EmoPair dataset, which leads the transition from the source emotion to the target emotion. To achieve this, we introduce an alignment loss in a neuro-symbolic space, between emotion embeddings $e$ from $\tau_{\theta}$ and text embeddings $c$ of instructions in each image pair of our EmoPair dataset with pre-trained CLIP. Specifically, the alignment loss is formulated as the inverse of cosine similarity between $e$ and $c$: $L_{emb}=1-\cos(e,c)$. Overall, we conduct joint training of $\tau_{\theta}$ and $\epsilon_{\theta}$ with the following loss: $L_{total}=L_{noise}+\lambda L_{emb}$, where $\lambda=0.5$ represents the weight for balancing different losses.
@@
In the first iteration of image generation, the source image $I_{s}$ serves as both the input and condition image, producing latent variables $z$ and $z_{t}$. These, along with the emotion vector $e$, are fed to the EmoEditor to obtain the generated image $\hat{I}$. The critic evaluates the generated image quality based on two criteria: (1) the structural similarity (SSIM) between $\hat{I}$ and $I_{s}$ falls within 0.3-0.8, and (2) the predicted emotion by $\mathcal{P}$ on $\hat{I}$ matches the target emotion, with a confidence level exceeding 0.8. If both criteria are met, EmoEditor ceases image generation; otherwise, it continues to generate new images. In subsequent iterations, we use $\hat{I}$ as the input image, generating its noisy latent variable $\hat{z}_{t}$, with $I_{s}$ remaining as the condition. Along with the emotion vector $e$, the EmoEditor generates new edited images for the critic. This process iterates until both criteria are satisfied, or the number of critic iterations exceeds a predefined limit of 30. In the latter case, we select the generated result with the highest confidence level predicted by $\mathcal{P}$ given the target emotion among all past iterations.
See Supp.Sec.S2.2 for Implementation Details.